Economic Order Quantity

Jane wants to determine the optimum amount of money to withdraw from an automatic teller machine (ATM) per transaction. The bank charges $.30 per ATM withdrawal transaction and a flat service charge of $5.00 per month. Jane spends an average of $10.00 per day. She figures there is a 10% chance that she will lose her wallet or be robbed in any given year. The bank pays 6% per year on checking account balances. (a) What is her optimal withdrawal amount per transaction? (b) How might the amount of Jane's withdrawals be altered if she moved to a high crime area?
Answer #1

Q.2

(a) Economic Order Quantity= √2DS/H

where: D=Demand in units(on annual basis), S = Order Cost(per purchase order). H = Holding Cost(per unit per year)

=√(2*38000*9)/1= √684000/1=827 units


answered by: ANURANJAN SARSAM